# Divide-Then-Aggregate: An Efficient Tool Learning Method via Parallel Tool Invocation

### v1

@@ Abstract
Although current Large Language Models (LLMs) exhibit impressive capabilities, performing complex real-world tasks still requires tool learning. Mainstream methods, such as CoT/ReAct, rely on step-by-step tool invocation to interact with external environments, but they are limited in perceptual scope and lack adequate task-planning capability. To address these limitations, other studies introduce the first Search-based Decision Tree (DFSDT), which still suffers from the high computational cost. In this paper, we introduce a novel parallel tool invocation paradigm, DTA-Llama (Divide-Then-Aggregate Llama). First, we transform traditional tree-based tool search paths into Directed Acyclic Graph (DAG) structure, generating a high-quality parallel tool invocation dataset. The DTA-Llama is then trained on the dataset to learn to iteratively divide the current task into several parallel tool invocation sub-tasks and aggregate the invocation results to decide the next actions. Furthermore, we introduce an efficient inference framework inspired by the Process/Threads mechanism when applying the DTA-Llama to practical tasks. Experimental results show that our approach substantially enhances task performance while reducing token consumption and inference time. Llama2-7B, using our method, is comparable to the official parallel function calling method of GPT-3.5. The relevant code, dataset, and model weights are available at https://corn0205.github.io/.
@@
Dongsheng Zhu 1, Weixian Shi 1, Zhengliang Shi 1,2 ‡, Zhaochun Ren 1, Shuaiqiang Wang 3, Lingyong Yan1, Dawei Yin 1 1 Baidu. Inc., Beijing, China 2 Shandong University, Qingdao, China 3 Leiden University, Leiden, The Netherlands {zhudongsheng, shiweixian, yanlingyong, yindawei}@baidu.com
@@ 1 Introduction
Large Language Models (LLMs), which are pre-trained and fine-tuned on massive amounts of textual data, have demonstrated powerful proficiency in various artificial intelligence tasks, such as conversation (Zheng et al., 2023), logical reasoning (Pan et al., 2023) and coding (Nijkamp et al., 2023). However, more real-world tasks often require the LLMs to interact with the environment to get necessary external information or feedback, such as checking real-time flight status (Guan et al., 2024) and complicate calculation for data analysis (Sun et al., 2024). To this end, tool learning has emerged recently which aims to equip the LLMs with external tools and teach them how to leverage the tools to accomplish real-world tasks.
@@
Previous tool learning methods typically work in pipelined and tree-based paradigms. Concretely, early studies usually perform the tool invocation in a pipeline, such as the Chain-of-Thought (CoT) reasoning (Kojima et al., 2022), and the ReAct mechanism (Yao et al., 2022). In these methods, tool learning agents (i.e. the LLMs) interact with the environment through a Thought-Action-Observation framework (as shown in Figure 1(a)). However, these methods usually focus on reflecting and planning based on local observations, rather than globally perceiving and planning the whole task solving paths. In contrast, the tree-based methods, such as ToolLLM (Qin et al., 2023) and Toolchain* (Zhuang et al., 2023), adopt tree-based algorithms, e.g., Depth First Search based Decision Tree (DFSDT), to perform the global planning of tool invocation. However, these methods still confront the inevitable backtracking mechanism that iteratively retries, which can usually significantly increase token consumption and inference time (as shown in Figure 1(d) for an example). Moreover, both two kinds of methods suffer from the limitation that LLMs invoke only one tool at each round during tool planning, which narrows their perceptual scope and necessitates more rounds of tool invocation, thereby reducing overall efficiency.
@@
In this paper, we propose a novel tool using model—DTA-Llama (Divide-Then-Aggregate Llama), which supports the parallel invocation of tools within each round. Specifically, we first transform traditional tree-based tool search paths into the Directed Acyclic Graph (DAG) structure with a level-order traversal progression (as shown in Figure 1(c)), enabling the parallelization of tools compared to previous methods. Leveraging the widely-used dataset ToolBench (Qin et al., 2023), we thus construct a high-quality parallel tool invocation dataset, DTA-Tool. Subsequently, we train the Llama model on DTA-Tool to develop DTA-Llama. Furthermore, we also develop a parallel tool invocation framework inspired by the Process/Threads mechanism (Silberschatz et al., 2006) for inference. Here, Process is responsible for planning tool invocation and dividing the parallelizable tools into multiple threads; Subsequently, Threads execute independently according to the plan. After execution, an intermediate state lock is created to aggregate the results from all threads. This approach allows our method to shorten the tool invocation path and improve the efficiency of LLM tool invocation.
@@
We evaluate our method on StableToolBench Guo et al. (2024), the comprehensive and reliable real-world tools benchmark, assessing its performance in terms of solvable pass rate (SoPR), solvable win rate (SoWR), and actual computational cost. Compared to other methods, we achieve better tool invocation performance with reduced computational cost, even reaching performance comparable to GPT-3.5 (OpenAI, 2022) function calling with just fine-tuning Llama2-7B (Touvron et al., 2023b). Furthermore, we fine-tuned different LLMs to validate the generalization ability of our approach across various models.
@@ 2 Related Work
The agent tool learning aims to expand LLMs capabilities by teaching LLMs to use external tools. Many early studies Patil et al. (2023); Tang et al. (2023); Huang et al. (2023) focus on laying the groundwork for datasets yet exhibit limited variety in tool usage. To bridge this gap, Qin et al. (2023) developed a more comprehensive multi-tool benchmark and proposed an advanced tool invocation method using Depth First Search-based Decision Tree (DFSDT). Building on this, Zhuang et al. (2023) employed A* search algorithm for pruning, while Kim et al. (2023) adopted a compiler-based approach to parallelize tool invocation, both of which improved efficiency to some extent. Meanwhile, Du et al. (2024) and Chen et al. (2024) controlled the stability of LLM tool invocation through self-reflection and Direct Preference Optimization (DPO, Rafailov et al., 2024), respectively. Despite these advances, these methods remain rooted in tree-based search paradigms, lacking a broader perspective on task planning. Additionally, recent works have begun exploring tool creation and integration with agents, opening new avenues for research (Qian et al., 2023; Hao et al., 2024; Schick et al., 2024; Hao et al., 2024; Yuan et al., 2023).
@@
Task planning capability is a crucial factor for the success of LLMs in problem-solving. Some methods attempt to decompose tasks into sub-goals and then plan for each sub-goal sequentially (Huang et al., 2022; Lu et al., 2024; Qian et al., 2024). HuggingGPT (Shen et al., 2024) utilizes the LLM as a controller, responsible for decomposing human-input queries into sub-tasks and ultimately generating a comprehensive response. Plan-and-Solve (Wang et al., 2023) employs a two-stage instruction prompting approach: “Let’s first devise a plan" and “Let’s carry out the plan". ProPrompt Singh et al. (2023) converts natural language descriptions of problems into coding tasks. In other studies, researchers seek to interleave task decomposition and sub-task planning, advancing them dynamically (Gao et al., 2023; Wu et al., 2023; Yao et al., 2024). The Chain of Thought (CoT) series of work (Kojima et al., 2022; Wei et al., 2022) guides LLMs in reasoning about complex problems by constructing trajectories. ReAct (Yao et al., 2022) alternates between reasoning (the thought process) and planning (the action steps). Reflection (Shinn et al., 2024) builds upon ReAct by introducing a mechanism for LLMs to reflect on previous failures.
@@ 3 Methodology
In this section, we describe: (1) the shortcoming analysis of previous methods (§ 3.1); (2) how the DTA-Llama be trained based on constructed parallel tool using data (§ 3.2); (3) how to implement an efficient, Process/Threads-based parallel framework during inference (§ 3.3). The last two parts work together to ensure the robustness and parallelism of tool invocation of DTA-Llama.
@@ 3.1 Background
As aforementioned, most recent tool invocation studies are tree-based, which are first developed by ToolLLM Qin et al. (2023). Comparing to the ReAct (Yao et al., 2022) (or CoT (Wei et al., 2022)) that invokes tools through pipelined tool interaction, the tree-based methods replace the serial tool usage with the Depth First Search-based Decision Tree (DFSDT) algorithm. Consequentially, these methods increase the fault tolerance of LLMs and improves task planning capabilities. However, this comes at the cost of increased time complexity. This is mainly because that DFSDT typically generates longer tool invocation sequences due to its backtracking mechanism, which involves multiple attempts at new nodes. While this improves task completion rates, it sacrifices execution efficiency (see Figure 1 for an example).
@@ 3.2 Divide-then-Aggregate Tool Invocation
To steer the LLMs with capabilities of parallel tool invocation, it is critical to construct the corresponding finetuning datasets. However, in most previous methods, e.g., vanilla CoT/ReAct methods or tree search-based algorithms like DFSDT, LLMs typically rely on invoking one tool at a time, which is not consistent with our setups. To this end, we utilize this type of data to transform it from a serial structure into a parallel structure.
As shown in Figure 2, we first collect the serial successful tool innovation path from original tool searching trees. Given the tree-like tool searching trajectories generated by the tree search-based algorithm, it is inevitable that the trajectories contain redundant or erroneous paths. Therefore, we take the successful node series spanning from the root to the successful leaf node as successful path $\mathcal{P}$, and retain the nodes in $\mathcal{P}$; all nodes in other paths are filtered. If a node in the failed paths is also part of the successful path, it is retained.
Next, we utilize a powerful LLM, to identify whether any tools in $\mathcal{P}$ can be executed in parallel. We choose GPT-4-turbo OpenAI (2023) to perform this task. If GPT-4 detects parallelizable tools, it establishes their relationships and organizes them into a Directed Acyclic Graph (DAG), represented as $\mathcal{G}$. The feasibility of parallel execution depends on input-output dependencies and logical causal relationships. For further details on transforming to the DAG structure using GPT-4, refer to Appendix A. For some data of $\mathcal{P}$ that cannot be parallelized, we do not discard them but instead retain their original structure. Finally, we construct the tool invocation mechanism by performing the level-order traversal on $\mathcal{G}$, allowing tools on the same level to be controllably invoked in parallel and aggregate. This implements the Divide-Then-Aggregate approach during the data construction phase.
It is important that we only transform the structure of $\mathcal{P}$ without modifying the semantic content. Therefore, the goal of our method is not data distillation.
@@
To ensure data quality, we apply filtering both before and after the data transformation. First, the raw serial data often contains incomplete tool invocation or task execution failures, which could negatively affect the LLMs fine-tuning. These cases are filtered out before the transformation. Additionally, even after the structural transformation, some structural errors may still persist. To address this issue, we have designed rule-based filtering. These rules leverage the acyclic nature of the DAG, along with its unique starting and ending points, to eliminate redundant tool invocation in cyclic invocation and situations where tool execution results cannot be aggregated. After these steps, we obtain DTA-Tool, a high-quality, DAG-based parallel tool invocation dataset with approximately 20k entries. We have documented the specific filtering rules in the appendix B.
@@
where $y$ represents the Thought generated by the LLMs, $q$ is the user instruction, and $o$ is the Observation. The Thought generated in the i-th round depends on the Thought and Observation from the previous round. In the final round, the focus of learning shifts from the Thought to the final answer generated by LLMs.
@@ 3.3 Process/Thread-based Inference
We fine-tuned the LLM with the DTA-Tool, resulting in DTA-Llama, which can invoke tools in parallel. To support the modified LLMs, we developed a new inference framework, as illustrated in the Figure 2(b). This framework redefines the Thought-Observation cycle based on CoT/ReAct, executing tool invocation in the form of Process/Threads.
@@
Originally, Thought could only design an invocation strategy for single tool, limiting its perceptual scope. In contrast, Process enhances the LLMs’ ability to divide tasks and plan multiple parallelizable tool invocation strategies. Specifically, during each round, LLMs first evaluate the task’s status and progress based on the historical trajectories. Then, LLMs analyze what needs to be done in the current step and decompose the task based on the available tools. Finally, LLMs sequentially generate a series of complete tool names along with their corresponding input parameters, which can be executed in parallel. This multi-tool approach helps broaden the perspective of LLMs, increasing the informational richness of each Thought step. In our framework, Process directly integrates Action into Thought. After careful deliberation, the LLMs provide a formalized tool invocation plan that can be extracted using regular expressions, facilitating subsequent execution by Threads.
@@
Threads refers to the steps faithfully execute the tool strategies presented in the Process. In previous frameworks, Thought provides only one tool invocation strategy. However, when Thought is capable of proposing parallel tool invocation strategies, the execution tool module must also be adapted accordingly. To address this, we introduced Threads. All the tool invocation strategies provided by the Process are distributed across multiple Threads, which then independently and concurrently execute each strategy. It is worth noting that real-time tool APIs generally allow a certain level of concurrency. Moreover, the tool invocation plans proposed by the Process are typically small in scale. Therefore, even if the parallel plans all target the same API, the concurrency is not high enough to cause an API crash.
@@
Table 2: A comparison different baselines and method on StableToolBench. Considering that real-world APIs are time-sensitive, the results of baselines presented in the table are reproduced during the period from September to October 2024 using their official implementation. We highlight the best performance of open-source models and GPT-series models with the red and blue, respectively.

### v2

@@ Abstract
While Large Language Models (LLMs) demonstrate remarkable capabilities, their ability to autonomously execute complex real-world tasks remains limited. Accordingly, tool learning has emerged to enable LLMs to effectively leverage external tools to extend their capabilities. Current tool-learning paradigms like CoT/ReAct employ sequential tool invocation but suffer from constrained perception and inadequate task planning. Alternative approaches using search-based decision trees incur substantial computational overhead. To address these limitations, we propose DTA-Llama (Divide-Then-Aggregate Llama), a novel parallel tool invocation framework featuring: (1) A Directed Acyclic Graph (DAG) structure that transformed from traditional tree-based tool search paths, enabling parallel execution and contributing high-quality training data; (2) A process-thread-inspired inference mechanism that iteratively decomposes tasks into parallel tool-using subtasks while aggregating results for subsequent decisions. Experimental results show that our approach substantially enhances task performance while reducing token consumption and inference time. Llama2-7B, using our method, is comparable to the official parallel function calling method of GPT-3.5. The relevant code, dataset, and model weights are available at https://corn0205.github.io/.
@@
Dongsheng Zhu1, Weixian Shi1, Zhengliang Shi2 ‡, Zhaochun Ren3, Shuaiqiang Wang1, Lingyong Yan1, Dawei Yin1 1Baidu Inc., Beijing, China   2Shandong University, Qingdao, China 3Leiden University, Leiden, The Netherlands {zhudongsheng, shiweixian, yanlingyong, yindawei}@baidu.com zhengliang.shii@gmail.com
@@ 1 Introduction
Large Language Models (LLMs), which are pre-trained and fine-tuned on massive amounts of textual data, have demonstrated powerful proficiency in various artificial intelligence tasks, such as conversation (Zheng et al., 2023; Zhu et al., 2024b), logical reasoning (Pan et al., 2023) and coding (Nijkamp et al., 2023). However, more real-world tasks often require the LLMs to interact with the environment to get necessary external information or feedback, such as checking real-time flight status (Guan et al., 2024) and complicate calculation for data analysis (Sun et al., 2024). To this end, tool learning has emerged recently which aims to equip the LLMs with external tools and teach them how to leverage the tools to accomplish real-world tasks.
@@
Previous tool learning methods typically work in pipelined and tree-based paradigms. Concretely, early studies usually perform the tool invocation in a pipeline, such as the Chain-of-Thought (CoT) reasoning (Kojima et al., 2022), and the ReAct mechanism (Yao et al., 2022). In these methods, tool learning agents (i.e., the LLMs) interact with the environment through a Thought-Action-Observation framework (as shown in Figure 1(a)). However, these methods usually focus on reflecting and planning based on local observations, rather than globally perceiving and planning the whole task solving paths. In contrast, the tree-based methods, such as ToolLLM (Qin et al., 2023) and Toolchain* (Zhuang et al., 2023), adopt tree-based algorithms, e.g., Depth First Search based Decision Tree (DFSDT), to perform the global planning of tool invocation. However, these methods still confront the inevitable backtracking mechanism that iteratively retries, which can usually significantly increase token consumption and inference time (as shown in Figure 1(d) for an example). Moreover, both two kinds of methods suffer from the limitation that LLMs invoke only one tool at each round during tool planning, which narrows their perceptual scope and necessitates more rounds of tool invocation, thereby reducing overall efficiency.
@@
In this paper, we introduce a novel model-based tool, DTA-Llama (Divide-Then-Aggregate Llama), which enables parallel tool invocation within each round of execution. Specifically, we convert traditional tree-based tool search paths into a Directed Acyclic Graph (DAG) structure with level-order traversal (as illustrated in Figure 1(c)), allowing for parallel execution of tools compared to previous sequential methods. Using the widely adopted ToolBench dataset (Qin et al., 2023), we construct a high-quality parallel tool invocation dataset, DTA-Tool, and train Llama (Touvron et al., 2023a, b; Dubey et al., 2024) on it to develop DTA-Llama. Additionally, we design a parallel tool invocation framework inspired by the Process/Threads mechanism (Silberschatz et al., 2006) for inference. In this framework, the Process component plans the tool invocation and divides parallelizable tools into separate Threads, which then execute independently according to the plan. After execution, an intermediate state lock aggregates the results from all threads. This design shortens the invocation path and significantly improves the efficiency of tool use in LLMs.
@@
We evaluate our approach on StableToolBench Guo et al. (2024), a comprehensive and reliable real-world tool-use benchmark. Performance is measured using solvable pass rate (SoPR), solvable win rate (SoWR), and actual computational cost. Compared to existing methods, our approach achieves superior tool invocation performance while reducing computational cost—even matching GPT-3.5’s (OpenAI, 2022) function-calling performance using only a fine-tuned Llama2-7B (Touvron et al., 2023b). To further assess generalization, we fine-tune multiple LLMs, demonstrating the robustness and generalization ability of our method across different models.
@@ 2 Related Work
The agent tool learning aims to expand LLMs capabilities by teaching LLMs to use external tools. Many early studies Patil et al. (2023); Tang et al. (2023); Huang et al. (2023) focus on laying the groundwork for datasets yet exhibit limited variety in tool usage. To bridge this gap, Qin et al. (2023) developed a more comprehensive multi-tool benchmark and proposed an advanced tool invocation method using Depth First Search-based Decision Tree (DFSDT). Building on this, Zhuang et al. (2023) employed A* search algorithm for pruning, while Kim et al. (2023) adopted a compiler-based approach to parallelize tool invocation, both of which improved efficiency to some extent. Meanwhile, Du et al. (2024) and Chen et al. (2024) controlled the stability of LLM tool invocation through self-reflection and Direct Preference Optimization (DPO, Rafailov et al., 2024), respectively. Despite these advances, these methods remain rooted in tree-based search paradigms, lacking a broader perspective on task planning. Additionally, recent works have begun exploring tool creation and integration with agents, opening new avenues for research (Qian et al., 2023; Yuan et al., 2023; Zhu et al., 2024a; Hao et al., 2024; Schick et al., 2024; Hao et al., 2024; Ma et al., 2024; Shen et al., 2024a; Yuan et al., 2024).
@@
Task planning capability is a crucial factor for the success of LLMs in problem-solving. Some methods attempt to decompose tasks into sub-goals and then plan for each sub-goal sequentially (Huang et al., 2022; Hu et al., 2023; Lu et al., 2024; Qian et al., 2024; Wang et al., 2024; Shi et al., 2024). HuggingGPT (Shen et al., 2024b) utilizes the LLM as a controller, responsible for decomposing human-input queries into sub-tasks and ultimately generating a comprehensive response. Plan-and-Solve (Wang et al., 2023) employs a two-stage instruction prompting approach: “Let’s first devise a plan" and “Let’s carry out the plan". ProPrompt Singh et al. (2023) converts natural language descriptions of problems into coding tasks. In other studies, researchers seek to interleave task decomposition and sub-task planning, advancing them dynamically (Gao et al., 2023; Wu et al., 2023; Yao et al., 2024; Shi et al., 2025). The Chain of Thought (CoT) (Kojima et al., 2022; Wei et al., 2022) guides LLMs in reasoning about complex problems by constructing trajectories. ReAct (Yao et al., 2022) alternates between reasoning (the thought process) and planning (the action steps). Reflection (Shinn et al., 2024) builds upon ReAct by introducing a mechanism for LLMs to reflect on previous failures.
@@ 3 Methodology
In this section, we describe: (1) the shortcoming analysis of previous methods (§ 3.1); (2) how the DTA-Llama be trained based on constructed parallel tool using data (§ 3.2); (3) how to implement an efficient, Process/Threads-based parallel framework during inference (§ 3.3).
@@ 3.1 Background
As aforementioned, most recent tool invocation studies are tree-based, which are first developed by ToolLLM Qin et al. (2023). Compared to the ReAct (Yao et al., 2022) (or CoT (Wei et al., 2022)) that invokes tools through pipelined tool interaction, the tree-based methods replace the serial tool usage with the Depth First Search-based Decision Tree (DFSDT) algorithm. Consequentially, these methods increase the fault tolerance of LLMs and improve task planning capabilities. However, this comes at the cost of increased time complexity. This is mainly because DFSDT typically generates longer tool invocation sequences due to its backtracking mechanism, which involves multiple attempts at new nodes. While this improves task completion rates, it sacrifices execution efficiency (see Figure 1 for an example).
@@ 3.2 Divide-then-Aggregate Tool Invocation
To steer the LLMs with the capabilities of parallel tool invocation, it is critical to construct the corresponding finetuning datasets. However, in most previous methods, e.g., vanilla CoT/ReAct methods or tree search-based algorithms like DFSDT, LLMs typically rely on invoking one tool at a time, which is not consistent with our setups. To this end, we utilize this type of data to transform it from a serial structure into a parallel structure.
As shown in Figure 2, we first collect the serial successful tool innovation path from original tool searching trees. Given the tree-like tool searching trajectories generated by the tree search-based algorithm, it is inevitable that the trajectories contain redundant or erroneous paths. Therefore, we define the node series spanning from the root to the successful leaf node as successful path $\mathcal{P}$; we retain only the nodes in $\mathcal{P}$, while filtering out other nodes.
Next, we utilize a powerful LLM, to identify whether any tools in $\mathcal{P}$ can be executed in parallel. We choose GPT-4-turbo OpenAI (2023) to perform this task. If GPT-4 detects parallelizable tools, it establishes their relationships and organizes them into a Directed Acyclic Graph (DAG), represented as $\mathcal{G}$. The feasibility of parallel execution depends on input-output dependencies and logical causal relationships. For the nodes that cannot be parallelized, we retain their original structure***For further details on transforming to the DAG structure using GPT-4, refer to Appendix A. .
Finally, we construct the tool invocation mechanism by performing the level-order traversal on $\mathcal{G}$, enabling tools at the same level to be executed in parallel in a controlled manner and their outputs to be aggregated accordingly. This process embodies the Divide-Then-Aggregate strategy during the data construction phase. It is worth noting that we only transform the structure of $\mathcal{P}$ without modifying the semantic content. Therefore, the objective of our approach is not to distill or compress the data, but rather to optimize its execution structure.
@@
To ensure data quality, we apply filtering both before and after the data transformation. On the one hand, we filter out the raw serial data that contain incomplete tool invocation or cause task execution failures. Because including these noisy data could negatively affect the subsequent LLMs fine-tuning. On the other hand, we devise a rule-based filtering method to reduce some structural errors after the structural transformation. These rules leverage the acyclic nature of the DAG, along with its unique starting and ending points, to eliminate redundant tool invocation in cyclic invocation and situations where tool execution results cannot be aggregated. After these steps, we obtain DTA-Tool, a high-quality, DAG-based parallel tool invocation dataset with approximately 20k entries. We have documented the specific filtering rules in the appendix B.
@@
where $y$ represents the Thought generated by the LLMs, $q$ is the user instruction, and $o$ is the Observation. The Thought generated in the i-th round depends on the Thought and Observation from the previous round. In the final round, the focus of learning shifts from the Thought to the final answer generated by LLMs. Finally, we fine-tuned the LLM with the DTA-Tool, resulting in DTA-Llama, which can invoke tools in parallel.
@@ 3.3 Process/Thread-based Inference
To support the modified LLMs, we developed a new inference framework, as illustrated in Figure 2(b). This framework redefines the Thought-Observation cycle based on CoT/ReAct, executing tool invocation in the form of Process/Threads.
@@
Originally, Thought could only design an invocation strategy for a single tool, limiting its perceptual scope. In contrast, Process enhances the LLMs’ ability to divide tasks and plan multiple parallelizable tool invocation strategies. Specifically, during each round, LLMs first evaluate the task’s status and progress based on the historical trajectories. Then, LLMs analyze what needs to be done in the current step and decompose the task based on the available tools. Finally, LLMs sequentially generate a series of complete tool names along with their corresponding input parameters, which can be executed in parallel. This multi-tool approach helps broaden the perspective of LLMs, increasing the informational richness of each Thought step. In our framework, Process directly integrates Action into Thought. After careful deliberation, the LLMs provide a formalized tool invocation plan that can be extracted using regular expressions, facilitating subsequent execution by Threads.
@@
Threads refers to the steps that faithfully execute the tool strategies presented in the Process. In previous frameworks, Thought provides only one tool invocation strategy. However, once Thought is capable of proposing multiple tool strategies in parallel, the execution component must also support concurrent processing. To this end, we introduce Threads. All the tool invocation strategies provided by the Process are distributed across multiple Threads, which then independently and concurrently execute each strategy. Importantly, most real-time tool APIs inherently support a moderate degree of concurrency. Furthermore, the tool invocation plans proposed by Process are typically lightweight. As a result, even when multiple concurrent invocations target the same API, the level of concurrency remains within a tolerable range and does not pose a risk of overloading the service.
@@
Table 2: A comparison different baselines and method on StableToolBench. Considering that real-world APIs are time-sensitive, the results of baselines presented in the table are reproduced during the period from September to October 2024 using their official implementation. We highlight the best performance of GPT-series models and open-source models with the red and blue, respectively.